1. Org Pvt. Ltd. is considering two mutually exclusive capital investments. The project’s expected net cash flows are as follows:

Expected Cash Flows Year Project A Project B

0 -400 -575

1 95 150

2 110 200

3 118 250

4 125 275

5 140 230

6 150 180

a. If you were told that each project’s cost of capital was 10%, which project should be selected using the NPV criteria? b. What is each project’s IRR? c. What is the regular payback period for these two projects? d. What is the profitability index for each project if the cost of capital is 12%? 


1
Expert's answer
2021-05-07T10:26:04-0400



(a) The net present value of project B is more, So I selected project B.


(b) The IRR for project A and B is 3% and 7%.


(c) Regular payback for these projects "A\text{ is }= 839-400=439 \text{ and for } B =1360.66-575=785.66"


(d) Profitability index at 12% fro Project "A=795.87-400=395.87" and for "B=1288.20-575=718.2"